Mass m1 = 5.65−kg is connected to mass m2 = 4.35−kg by a string that passes through a massless and frictionless pulley. Mass m2 is connected to mass m3 = 3.00−kg by a string. find t1 and t2 in N
Answer #1

m. = 5.65kg, m2 = 4.35kg, m3=3.00kg - ti ti mal mig um. tz m3 I m3 g for mi ti-mig for m2 =m m, a 22-t, + m₂g=mza-@= 1 for m₂ mag-tz = mza - tz M3 g-m₂ a in mzg-mza-t, tm zg=ma g(m₂ tmz)-t, = a (Mat m3) g(4.35+3) - t;= a(4.35+3) 7.35 7-6, tin @ t₂ = m₂lg-a)=3(9.8-1-7X9.8) 13 tz = 25.55N